1.The following table gives mean daily calorie intake (all residents) and infant mortality rate (per 1000 births) for ten countries.

Country                Mean Daily Calories           Infant Mortality Rate

     Afghanistan                      1523                              154

             Austria                         3495                                 6

             Burundi                        1941                              114

            Colombia                       2678                               24

             Ethiopia                         1610                             107

            Germany                        3443                                6

              Liberia                          1640                             153

         New Zealand                     3362                                 7

             Turkey                          3429                               44

               USA                           3671                                 7

a)      Find the regression equation.

b)   Calculate the correlation coefficient, accurate to three decimal places  

c) Test the significance of the correlation coefficient with a 0.05 level of significance. Include definition of hypotheses, test statistic, p-value or critical value, decision about Ho, and conclusion.

d)      Find the best estimate for the infant mortality rate for a country with mean daily calorie intake of 2800.

PLEASE SHOW ALL THE WORK AND FOLLOW DIRECTIONS ACCURATE AND MARK EACH PART A), B), C), D)

THANK YOU
Answer #1
x y (x-x̅)² (y-ȳ)² (x-x̅)(y-ȳ)
1523 154 1336798.4400 8427.2400 -106139.160
3495 6 665529.6400 3158.4400 -45847.960
1941 114 544939.2400 2683.2400 -38238.760
2678 24 1.4400 1459.2400 45.840
1610 107 1143188.6400 2007.0400 -47900.160
3443 6 583390.4400 3158.4400 -42925.560
1640 153 1079936.6400 8244.6400 -94359.360
3362 7 466215.8400 3047.0400 -37690.560
3429 44 562200.0400 331.2400 -13646.360
3671 7 983667.2400 3047.0400 -54747.360
ΣX ΣY Σ(x-x̅)² Σ(y-ȳ)² Σ(x-x̅)(y-ȳ)
total sum 26792.00 622.00 7365867.60 35563.6 -481449.4
mean 2679.20 62.20 SSxx SSyy SSxy

a)

Sample size,   n =   10      
here, x̅ = Σx / n=   2679.200          
ȳ = Σy/n =   62.200          
SSxx =    Σ(x-x̅)² =    7365867.6000      
SSxy=   Σ(x-x̅)(y-ȳ) =   -481449.4      
              
estimated slope , ß1 = SSxy/SSxx =   -481449.4/7365867.6=   -0.0654      
intercept,ß0 = y̅-ß1* x̄ =   62.2- (-0.0654 )*2679.2=   237.3184      
              
Regression line is, Ŷ=   237.318   + (   -0.065   )*x

b) correlation coefficient ,    r = SSxy/√(SSx.SSy) =   -0.941

c)

Ho:   ρ = 0  
Ha:   ρ ╪ 0  
n=   10  
alpha,α =    0.05  
correlation , r=   -0.9407  
t-test statistic = r*√(n-2)/√(1-r²) =        -7.8407
DF=n-2 =   8  
p-value =    0.0001  
Decison:   p value < α , So, Reject Ho  


critical t-value =    2.3060  
decision: reject Ho, because test stat > critical value

There is enough evidence of significant correlation

d) Predicted Y at X=   2800   is          
Ŷ=   237.31844   +   -0.06536   *2800=   54.30

please revert for doubts and
please UPVOTE the solution